# Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples

### v1

@@ 3.1 Maximum Eccentricity
To assess the role of a tertiary companion in our CHE binary at the ZAMS, we first focus on the maximum eccentricity that can be induced via the ZLK mechanism.
Figure 5 shows the maximum eccentricity for 8 different models (see §2.2 for details). The left column of Fig. 5 shows the DA approximation, which neglects perturbations occurring on timescales shorter than the outer orbital period (e.g., Mangipudi et al., 2022), while the right column includes models with the SA approximation. From top to bottom, models account for short-range forces:
@@
  • Panel (e) shows the results when accounting for the equilibrium tide effect on the inner binary, where tidal bulges raised on the stellar surfaces induce a quadrupole moment that leads to apsidal precession (e.g., Hamers et al., 2021, and references therein). Tides efficiently suppress the ZLK mechanism, leading to only small eccentricities ($e_{\rm max}\lesssim 0.2$) when the outer orbital period is long ($P_{\rm out}\gtrsim 100$ d), spefically when $\epsilon_{\rm Tide}\gg 1$. Even for the most massive ($m_{3}=100M_{\odot}$) and closest ($P_{\rm out}\approx 5$ d) tertiaries in our parameter space, the combined effect of ZLK and tides only leads to moderately large eccentricities ($e_{\rm max}\lesssim 0.75$).
@@ 3.2 Fraction of Stellar Mergers
In §3.1, we utilised our default CHE binary model at the ZAMS (§2.1.1) to determine the maximum eccentricity attained in a triple configuration with a tertiary mass in the range $1\leq m_{3}/M_{\odot}\leq 100$ and outer orbital period $1\leq P_{\rm{out}}/\rm{d}\leq 1000$. Those results represent a single realisation within the parameter space of initial mutual inclinations, corresponding to the angle in our grid that results in the largest eccentricity. In this section, we examine the whole range of inclinations to assess the fraction of stellar mergers (see §2.2.3 for our criteria). For each point in the grid of $m_{3}$ and $a_{\rm out}$, we estimate a merger fraction ($f_{\rm merger}$). This fraction quantifies the proportion of inclinations, relative to a uniform distribution in $\cos(i_{0})$, that result in mergers.
@@
Figure 6: Merger analysis for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$), specifically addressing the von Zeipel-Lidov-Kozai mechanism including the single-averaged approximation, as well as the role of apsidal precession from general relativity and tides (§2.2.2). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. Specifically, we show the merger fraction ($f_{\rm merger}$, left panels), derived from an uniform grid in initial mutual inclinations inclinations ($\cos(i_{0})=\{-1,1\}$), as well as the value of $\cos(i_{0})$ that results in the maximum eccentricity (right panels). We display the analysis for our chemically homogenously evolving binary at the zero-age main sequence (ZAMS, top panels) and at $5.12$ Myr (bottom panels), shortly after the end of the main sequence (cf. Fig. 1). In the right panels, grey indicates the region where the torque induced by the tertiary on the inner binary is not able to overcome the effect of tides.
@@
In the top-right panel of Fig. 6, we present the value of $\cos(i_{0})|_{e_{\rm max}}$ at the ZAMS. We find that massive tertiaries in wide orbits have $\cos(i_{0})|_{e_{\rm max}}\approx\cos(90\ \rm{deg})=0$, which confirms the results in the point-mass test-particle approximation. However, when the test-particle approximation is relaxed, less massive tertiaries in tight orbits deviate from this value, and the maximum eccentricity is achieved in highly-inclined retrograde orbits. For low-mass ($m_{3}\lesssim 5M_{\odot}$) close ($P_{\rm out}\lesssim 50$ d) tertiaries, there exists a significant portion of the parameter space where retrograde orbits, characterized by $\cos(i_{0})|_{e_{\rm max}}=\cos(180\ \rm{deg})=-1$, lead to the maximum eccentricity. The transition from $\cos(i_{0})|_{e_{\rm max}}=0$ to $\cos(i_{0})|_{e_{\rm max}}=-1$ signifies not only a departure from the test-particle approximation ($\eta>1$), but also a regime where short-range forces start to play a role ($\epsilon_{\rm GR,Tide}>1$). In this regime, the ‘window’ of inclinations becomes narrower and more asymmetric around $\cos(i_{0})|_{e_{\rm max}}$ (cf. Fig. 1 in Anderson et al., 2017), resulting in a decreased number of mergers.
@@
These results underscore how stellar structure and binary evolution influence the dynamics of massive triple systems. We specifically highlight two regimes that likely occur promptly around the ZAMS or after the end of the main sequence, resulting in hydrogen-rich and helium-rich stellar mergers, respectively. We also show how the inclination that leads to the maximum eccentricity, along with the range of inclinations resulting in a threshold eccentricity required for merger, depends on the tertiary. For tertiary objects with masses $m_{3}\gtrsim 10M_{\odot}$, where the outer orbital angular momentum dominates ($\eta\ll 1$, Fig. 2) close to the test-particle limit, the inclination that leads to the maximum eccentricity is 90 deg. Conversely, for tertiary objects with $m_{3}<10M_{\odot}$, the inclination that results in maximum eccentricity is more sensitive to the stellar structure and may even be achieved in coplanar retrograde orbits.
Figure 7: Different outcomes in the parameter space of tight triples, with arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$), that host inner binaries whose component stars undergo chemically homogeneous evolution. At the zero-age main sequence, the system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. Grey and light grey represents the region where von Zeipel-Kozai-Lidov (ZLK) oscillations can lead to prompt hydrogen-rich and helium-rich stellar mergers in the inner binary, respectively. The colour bar gradient illustrates how ZLK oscillations can decrease the time-to-coalescence via gravitational-wave emission ($\tau_{\rm{merge,min}}$), potentially reducing it to the ZLK timescale (cf. Fig. 4). Wide triples ($P_{\rm{out}}\gtrsim 1000$ d) remain unperturbed, with the inner binary evolving into a binary black hole that can merge in approximately $450$ Myr (§2.1.1). The red dashed vertical line marks the mass where $m_{3}=m_{1,\rm{ZAMS}}$, roughly delineating systems where the tertiary may evolve more rapidly than either of the inner binary components, potentially leading to a tertiary mass-transfer episode.
@@ 3.3 Effects on Populations
After assessing the triple configurations that lead to stellar mergers (§3.2), in this section we examine the impact of a tertiary companion at various evolutionary stages of our standard CHE binary model (§2.1.1).
Figure 7 shows the summary of our analysis. We use different colours to highlight various regions of interest. The dark grey region indicates dynamically unstable triples. The grey region delineates the parameter space where a tertiary object can induce a hydrogen-rich merger between the inner stars, provided that the mutual inclination promotes the ZLK mechanism. This hydrogen-rich merger occurs near the ZAMS and on a timescale comparable to the ZLK timescale (Fig. 4). Similarly, the light grey region indicates the parameter space where an inner binary, which does not merge during the main sequence, can merge shortly afterward, resulting in a helium-rich stellar merger. This helium-rich merger would take place after the main sequence, either during the star’s rapid contraction phase or after this contraction has ceased. The parameter space of these mergers is narrower compared to hydrogen-rich mergers, yet it can lead to an intriguing merger remnant (see §4.1.1 for further discussion).
If a stellar merger is avoided, our default CHE binary can form a BBH. For simplicity, we consider the case in which a BBH forms with identical parameters to those at the end of the stellar model simulation (§2.1.1 and Fig. 1). Also in Fig. 7, we denote in colour the minimum time to coalescence via gravitational-wave emission ($\tau_{\rm merge,min}$) for the inner BBH. If the BBH evolves in isolation or unperturbed by the tertiary, the time-to-coalescence is $\tau_{\rm merge}\approx 450$ Myr. This is the case for triples with $P_{\rm out}>500$ d. However, there is a region of interest, dependant on tertiary mass, between $15\lesssim P_{\rm out}/\rm{d}\lesssim 500$. In this region, our default CHE binary might avoid a stellar merger and become a BBH. Once the BBH forms, tidal forces are no longer present (i.e., $\epsilon_{\rm Tide}=0$). This absence allows the maximum induced eccentricity via the ZLK mechanism to increase, now primarily quenched through GR effects. This induced eccentricity can lead to coalescence through gravitational-wave emission. As demonstrated by Liu & Lai (2018), the merger timescale of such BBH is given by
@@
Finally, in Fig. 7 we indicate with a red vertical dashed line where $m_{3}=m_{1,\rm{ZAMS}}$. In a coeval triple system, where all component stars are born at the same time, tertiaries with masses $m_{3}>m_{1,\rm{ZAMS}}$ will evolve more rapidly than the stars of the inner binary. This may lead to a tertiary mass-transfer episode, potentially influencing the future evolution of the triple system and causing deviations from the scenarios described here (e.g., de Vries et al., 2014).
Figure 8 shows the same exploration as in Fig. 7 but for a stellar model at $Z=0.0042$, which is twice as high as the commonly adopted value for the Small Magellanic Cloud (Sharpe et al., 2024). Except for the initial metallicity, which is ten times larger than in our default CHE (§2.1.1), all other initial parameters are identical (see App. A for details). At this higher metallicity, the main difference is the increased mass loss due to stellar winds, which remove a combined total of $\approx 57\ M_{\odot}$($\approx 16\ M_{\odot}$) during(after) the main sequence and expand the orbit by a factor of a few (App. A). This increased mass loss and orbital widening can lead to the formation of a BBH that would not merge within the age of the Universe. However, in the (tertiary mass-dependant) region where $P_{\rm out}\gtrsim 20$ d, there are many configurations where the ZLK mechanism can assist a merger within the age of the Universe, on timescales as short as the ZLK timescale. Overall, we find that ZLK mechanism in CHE binaries can result in prompt stellar or BBH mergers.
Figure 8: Same as Fig. 7 but for $Z=0.0042$. Values of $\tau_{\rm merge,min}<1$ yr are set to 1 year. The stellar model and merger analysis at this metallicity are provided in App. A.
@@ 4.1.1 Stellar Mergers
One of our goals in this paper is to establish a concise framework in order to investigate which orbital configurations lead to a stellar merger of the inner binary. These dynamical mergers occur rapidly (Vigna-Gómez et al., 2022), on timescales similar to the ZLK timescale, and reduce the triple system to a binary. Theory (e.g., Schneider et al., 2019) and observations (e.g., Frost et al., 2024) tentatively suggest that merger remnants of massive stars are highly magnetic, at least immediately after the merger, with additional predictions for slow rotation (e.g., Schneider et al., 2019; Wang et al., 2022; Keszthelyi et al., 2022) and unusual abundances due to enhanced mixing (e.g., Maeder & Meynet, 2005). Most merger scenarios discussed in the literature involve collisions (e.g., Hills & Day, 1976) or are initiated in isolated binary systems due to the radial expansion of at least one component star (e.g., Podsiadlowski et al., 1992), even when they occur within triples (Stegmann et al., 2022) or quadruples (Vigna-Gómez et al., 2022). Here, we explore a scenario in which triple dynamics can aid a stellar merger during two specific phases in the evolution of stars.
@@
Second, we examined helium-rich mergers after the main-sequence. This second phase occurs when CHE stars contract on a thermal timescale (Fig. 1). This contraction reduces the influence of tides, thereby diminishing the suppression of the ZLK mechanism. As a result, some binaries that did not merge during the main sequence may do so shortly thereafter. These helium-rich mergers were first identified in Dorozsmai et al. (2024) through a population analysis. Although the parameter space leading to helium-rich mergers is more restricted compared to hydrogen-rich mergers (Fig. 7 and 8), helium-rich mergers can produce unique remnants. In principle, they could result in massive highly-magnetised helium stars with $\lesssim 10^{5}$ yr remaining before stellar collapse. Such stellar structures have been previously linked to long-duration gamma-ray bursts and Type Ic superluminous supernovae (Yoon et al., 2006; Aguilera-Dena et al., 2018). Depending on the mass, they could also lead to (pulsational) pair-instability supernovae (e.g., Heger et al., 2003).
Finally, we highlighted the region within our parameter space that is likely to be influenced by the tertiary, assuming it is a star. This region is shown in Fig. 7 and 8 as a red dashed vertical line, indicating where $m_{3}=m_{1,\rm{ZAMS}}$. In coeval star systems, more massive stars evolve more rapidly than their less massive counterparts, causing them to leave the main sequence first. Canonical stellar models for non-rotating stars show rapid expansion shortly after the main sequence. In our tight triple system, such expansion could lead to tertiary tides (Gao et al., 2023) and eventually result in a mass transfer episode from the tertiary to the inner binary. The evolution and outcome of this tertiary mass transfer episode are quite uncertain (de Vries et al., 2014; Di Stefano, 2020b, a; Comerford & Izzard, 2020; Glanz & Perets, 2021a; Hamers et al., 2022). One possible outcome is a stellar merger of the inner binary (Dorozsmai et al., 2024), or the formation of very tight systems following a triple common-envelope episode (Glanz & Perets, 2021b; Vigna-Gómez et al., 2021). In any case, mass transfer from the tertiary will envelop the inner binary in a gaseous medium (e.g., Antoni et al., 2019), a configuration likely to affect the three-body dynamics.

### v2

@@ 3.1 Maximum eccentricity
To assess the role of a tertiary companion in our CHE binary at the ZAMS, we first focus on the maximum eccentricity that can be induced via the ZLK mechanism. Figure 5 shows the maximum eccentricity for 8 different models (see Sect. 2.2 for details). The left column of Fig. 5 shows the DA approximation, which neglects perturbations occurring on timescales shorter than the outer orbital period (e.g. Mangipudi et al., 2022), while the right column includes models with the SA approximation. From top to bottom, the models account for short-range forces, as described below.
@@
  • Panel (e) shows the results when accounting for the equilibrium tide effect on the inner binary, where tidal bulges raised on the stellar surfaces induce a quadrupole moment that leads to apsidal precession (e.g. Hamers et al., 2021, and references therein). Tides efficiently suppress the ZLK mechanism, leading to only small eccentricities ($e_{\rm max}\lesssim 0.2$) when the outer orbital period is long ($P_{\rm out}\gtrsim 100$ d), specifically when $\epsilon_{\rm Tide}\gg 1$. Even for the most massive ($m_{3}=100M_{\odot}$) and closest ($P_{\rm out}\approx 5$ d) tertiaries in our parameter space, the combined effect of ZLK and tides only leads to moderately large eccentricities ($e_{\rm max}\lesssim 0.75$).
@@ 3.2 Fraction of stellar mergers
In Sect. 3.1 we utilise our default CHE binary model at the ZAMS (Sect. 2.1.1) to determine the maximum eccentricity attained in a triple configuration with a tertiary mass in the range $1\leq m_{3}/M_{\odot}\leq 100$ and outer orbital period $1\leq P_{\rm{out}}/\rm{d}\leq 1000$. Those results represent a single realisation within the parameter space of initial mutual inclinations, corresponding to the angle in our grid that results in the largest eccentricity. In this section, we examine the whole range of inclinations to assess the fraction of stellar mergers (see Sect. 2.2.3 for our criteria). For each point in the grid of $m_{3}$ and $a_{\rm out}$, we estimate a merger fraction ($f_{\rm merger}$). This fraction quantifies the proportion of inclinations, relative to a uniform distribution in $\cos(i_{0})$, that result in mergers.
@@
Figure 6: Merger analysis for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$), specifically addressing the ZLK mechanism including the single-averaged approximation, as well as the role of apsidal precession from GR and tides (Sect. 2.2.2). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. Specifically, we show the merger fraction ($f_{\rm merger}$, left panels), derived from an uniform grid in initial mutual inclinations inclinations ($\cos(i_{0})=\{-1,1\}$), as well as the value of $\cos(i_{0})$ that results in the maximum eccentricity (right panels). We display the analysis for our chemically homogenously evolving binary at the ZAMS (top panels) and at $5.12$ Myr (bottom panels), shortly after the end of the main sequence (cf. Fig. 1). In the right panels, grey indicates the region where the torque induced by the tertiary on the inner binary is not able to overcome the effect of tides.
@@
In the top-right panel of Fig. 6, we present the value of $\cos(i_{0})|_{e_{\rm max}}$ at the ZAMS. We find that massive tertiaries in wide orbits have $\cos(i_{0})|_{e_{\rm max}}\approx\cos(90\ \rm{deg})=0$, which confirms the results in the point-mass test-particle approximation. However, when the test-particle approximation is relaxed, less massive tertiaries in tight orbits deviate from this value, and the maximum eccentricity is achieved in highly-inclined retrograde orbits. For low-mass ($m_{3}\lesssim 5M_{\odot}$) close ($P_{\rm out}\lesssim 50$ d) tertiaries, there exists a significant portion of the parameter space where retrograde orbits, characterised by $\cos(i_{0})|_{e_{\rm max}}=\cos(180\ \rm{deg})=-1$, lead to the maximum eccentricity. The transition from $\cos(i_{0})|_{e_{\rm max}}=0$ to $\cos(i_{0})|_{e_{\rm max}}=-1$ signifies not only a departure from the test-particle approximation ($\eta>1$), but also a regime where short-range forces start to play a role ($\epsilon_{\rm GR,Tide}>1$). In this regime, the ‘window’ of inclinations becomes narrower and more asymmetric around $\cos(i_{0})|_{e_{\rm max}}$ (cf. Fig. 1 in Anderson et al., 2017), resulting in a decreased number of mergers.
@@
These results underscore how stellar structure and binary evolution influence the dynamics of massive triple systems. We specifically highlight two regimes that likely occur promptly around the ZAMS or after the end of the main sequence, resulting in hydrogen-rich and helium-rich stellar mergers, respectively. We also show how the inclination that leads to the maximum eccentricity, along with the range of inclinations resulting in a threshold eccentricity required for merger, depends on the tertiary. For tertiary objects with masses $m_{3}\gtrsim 10M_{\odot}$, where the outer orbital angular momentum dominates ($\eta\ll 1$; Fig. 2) close to the test-particle limit, the inclination that leads to the maximum eccentricity is 90 deg. Conversely, for tertiary objects with $m_{3}<10M_{\odot}$, the inclination that results in maximum eccentricity is more sensitive to the stellar structure and may even be achieved in coplanar retrograde orbits.
Figure 7: Different outcomes in the parameter space of tight triples, with arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$), that host inner binaries whose component stars undergo CHE. At the ZAMS, the system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity of $Z=0.00042$. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. Grey and light grey represents the region where von Zeipel-Kozai-Lidov (ZLK) oscillations can lead to prompt hydrogen-rich and helium-rich stellar mergers in the inner binary, respectively. The colour bar gradient illustrates how ZLK oscillations can decrease the time-to-coalescence via gravitational-wave emission ($\tau_{\rm{merge,min}}$), potentially reducing it to the ZLK timescale (cf. Fig. 4). Wide triples ($P_{\rm{out}}\gtrsim 1000$ d) remain unperturbed, with the inner binary evolving into a BBH that can merge in approximately $450$ Myr (Sect. 2.1.1). The vertical dashed red line marks the mass where $m_{3}=m_{1,\rm{ZAMS}}$, roughly delineating systems where the tertiary may evolve more rapidly than either of the inner binary components, potentially leading to a tertiary mass-transfer episode.
@@ 3.3 Effects on populations
After assessing the triple configurations that lead to stellar mergers (Sect. 3.2), in this section we examine the impact of a tertiary companion at various evolutionary stages of our standard CHE binary model (Sect. 2.1.1). Figure 7 shows the summary of our analysis. We use different colours to highlight various regions of interest. The dark grey region indicates dynamically unstable triples. The grey region delineates the parameter space where a tertiary object can induce a hydrogen-rich merger between the inner stars, provided that the mutual inclination promotes the ZLK mechanism. This hydrogen-rich merger occurs near the ZAMS and on a timescale comparable to the ZLK timescale (Fig. 4). Similarly, the light grey region indicates the parameter space where an inner binary, which does not merge during the main sequence, can merge shortly afterwards, resulting in a helium-rich stellar merger. This helium-rich merger would take place after the main sequence, either during the star’s rapid contraction phase or after this contraction has ceased. The parameter space of these mergers is narrower compared to hydrogen-rich mergers, yet it can lead to an intriguing merger remnant (see Sect. 4.1.1 for further discussion).
If a stellar merger is avoided, our default CHE binary can form a BBH. For simplicity, we consider the case in which a BBH forms with identical parameters to those at the end of the stellar model simulation (Sect. 2.1.1 and Fig. 1). Also in Fig. 7, we denote in colour the minimum time to coalescence via gravitational-wave emission ($\tau_{\rm merge,min}$) for the inner BBH. If the BBH evolves in isolation or unperturbed by the tertiary, the time-to-coalescence is $\tau_{\rm merge}\approx 450$ Myr. This is the case for triples with $P_{\rm out}>500$ d. However, there is a region of interest, dependant on tertiary mass, between $15\lesssim P_{\rm out}/\rm{d}\lesssim 500$. In this region, our default CHE binary might avoid a stellar merger and become a BBH. Once the BBH forms, tidal forces are no longer present (i.e. $\epsilon_{\rm Tide}=0$). This absence allows the maximum induced eccentricity via the ZLK mechanism to increase, now primarily quenched through GR effects. This induced eccentricity can lead to coalescence through gravitational-wave emission. As demonstrated by Liu & Lai (2018), the merger timescale of such BBH is given by
@@
Finally, in Fig. 7 we indicate with a vertical dashed red line where $m_{3}=m_{1,\rm{ZAMS}}$. In a coeval triple system, where all component stars are born at the same time, tertiaries with masses $m_{3}>m_{1,\rm{ZAMS}}$ will evolve more rapidly than the stars of the inner binary. This may lead to a tertiary mass-transfer episode, potentially influencing the future evolution of the triple system and causing deviations from the scenarios described here (e.g. de Vries et al., 2014).
Figure 8 shows the same exploration as in Fig. 7 but for a stellar model at $Z=0.0042$, which is twice as high as the commonly adopted value for the Small Magellanic Cloud (SMC; Sharpe et al., 2024). Except for the initial metallicity, which is ten times larger than in our default CHE (Sect. 2.1.1), all other initial parameters are identical (see Appendix A for details). At this higher metallicity, the main difference is the increased mass loss due to stellar winds, which remove a combined total of $\approx 57\ M_{\odot}$($\approx 16\ M_{\odot}$) during(after) the main sequence and expand the orbit by a factor of a few (Appendix A). This increased mass loss and orbital widening can lead to the formation of a BBH that would not merge within the age of the Universe. However, in the (tertiary mass-dependant) region where $P_{\rm out}\gtrsim 20$ d, there are many configurations where the ZLK mechanism can assist a merger within the age of the Universe, on timescales as short as the ZLK timescale. Overall, we find that ZLK mechanism in CHE binaries can result in prompt stellar or BBH mergers.
Figure 8: Same as Fig. 7 but for $Z=0.0042$. Values of $\tau_{\rm merge,min}<1$ yr are set to 1 year. The stellar model and merger analysis at this metallicity are provided in Appendix A.
@@ 4.1.1 Stellar mergers
One of our goals in this paper is to establish a concise framework in order to investigate which orbital configurations lead to a stellar merger of the inner binary. These dynamical mergers occur rapidly (Vigna-Gómez et al., 2022), on timescales similar to the ZLK timescale, and reduce the triple system to a binary. Theory (e.g. Schneider et al., 2019) and observations (e.g. Frost et al., 2024) tentatively suggest that merger remnants of massive stars are highly magnetic, at least immediately after the merger, with additional predictions for slow rotation (e.g. Schneider et al., 2019; Wang et al., 2022; Keszthelyi et al., 2022) and unusual abundances due to enhanced mixing (e.g. Maeder & Meynet, 2005). Most merger scenarios discussed in the literature involve collisions (e.g. Hills & Day, 1976) or are initiated in isolated binary systems due to the radial expansion of at least one component star (e.g. Podsiadlowski et al., 1992), even when they occur within triples (Stegmann et al., 2022) or quadruples (Vigna-Gómez et al., 2022). Here, we explore a scenario in which triple dynamics can aid a stellar merger during two specific phases in the evolution of stars.
@@
Second, we examined helium-rich mergers after the main-sequence. This second phase occurs when CHE stars contract on a thermal timescale (Fig. 1). This contraction reduces the influence of tides, thereby diminishing the suppression of the ZLK mechanism. As a result, some binaries that did not merge during the main sequence may do so shortly thereafter. These helium-rich mergers were first identified in Dorozsmai et al. (2024) through a population analysis. Although the parameter space leading to helium-rich mergers is more restricted compared to hydrogen-rich mergers (Fig. 7 and 8), helium-rich mergers can produce unique remnants. In principle, they could result in massive highly magnetised helium stars with $\lesssim 10^{5}$ yr remaining before stellar collapse. Such stellar structures have been previously linked to long-duration gamma-ray bursts and Type Ic superluminous supernovae (Yoon et al., 2006; Aguilera-Dena et al., 2018). Depending on the mass, they could also lead to (pulsational) pair-instability supernovae (e.g. Heger et al., 2003).
Finally, we highlighted the region within our parameter space that is likely to be influenced by the tertiary, assuming it is a star. This region is shown in Fig. 7 and 8 as a vertical dashed red line, indicating where $m_{3}=m_{1,\rm{ZAMS}}$. In coeval star systems, more massive stars evolve more rapidly than their less massive counterparts, causing them to leave the main sequence first. Canonical stellar models for non-rotating stars show rapid expansion shortly after the main sequence. In our tight triple system, such expansion could lead to tertiary tides (Gao et al., 2023) and eventually result in a mass transfer episode from the tertiary to the inner binary. The evolution and outcome of this tertiary mass transfer episode are quite uncertain (de Vries et al., 2014; Di Stefano, 2020b, a; Comerford & Izzard, 2020; Glanz & Perets, 2021a; Hamers et al., 2022). One possible outcome is a stellar merger of the inner binary (Dorozsmai et al., 2024), or the formation of very tight systems following a triple common-envelope episode (Glanz & Perets, 2021b; Vigna-Gómez et al., 2021). In any case, mass transfer from the tertiary will envelop the inner binary in a gaseous medium (e.g. Antoni et al., 2019), a configuration likely to affect the three-body dynamics (see also Shariat et al., 2025b; Sciarini et al., 2025, for other recent investigations on the dynamics of intermediate-mass and high-mass triples, respectively).